### v4

@@ VII Related Work
Code generation tasks is about creating source code from provided natural language descriptions or requirements, which has been a longstanding focal point in software engineering research.
@@
Retrieval Augmented Code Generation. Retrieval Augmented Code Generation (RACG)Ā [65, 62, 66, 67] represents an emerging and promising area of research that combines the strengths of retrieval-based methods with code generation techniques. Hayati et al.Ā [62] introduced the RECODE model, which is the first to incorporate information retrieval techniques into the code generation task. Parvez et al.Ā [66] proposed the REDCODER, which utilizes retrieval to enhance the input of the generation model, aiming to maximize the utilization of additional information. Drain et al.Ā [65] retrieve template functions from GitHub and relevant answers from Stack Overflow to improve code generation. Zhou et al.Ā [67] proposed Docprompting, which retrieve relevant APIs from documents and add them into prompts to improve code generation. Zan et al.Ā [68] proposed CodeGenAPI, which retrieve private APIs from documents to improve code generation. the APIRetriever module of CodeGenAPI and our approach both focus on private API recommendation. However, CodeGenAPI directly retrieves APIs from documentation and cannot effectively handle high-level tasks that require combinations of APIs.
@@ VIII Conclusion and Future Work
Code generation within code repositories often necessitates comprehensive knowledge of the entire repository context. Existing techniques often lack such a holistic understanding of code repository knowledge, which is crucial for effective code generation. In this paper, we introduce the $A^{3}$-CodGen framework, which is designed to construct, mine, and unify three essential types of code repository knowledge: local-aware, global-aware, and third-party-library-aware knowledge. To validate the effectiveness of these knowledge types, we conducted benchmark experiments and identified optimal configurations for their utilization. The results demonstrate that the $A^{3}$-CodGen framework significantly enhances the ability of Language Model Models (LLMs) to fully utilize local, global, and third-party library information. This results in the generation of concise, high-quality code. In the future, we plan to expand the applicability of our framework to multiple programming languages. Furthermore, we aim to enhance our knowledge utilization methods to enable LLMs to become more aware of various types of code knowledge within code repositories.

### v5

@@ VII Related Work
Code generation tasks is about creating source code from provided natural language descriptions or requirements, which has been a longstanding focal point in SE research.
@@
Retrieval Augmented Code Generation. Retrieval Augmented Code Generation (RACG)Ā [86, 82, 87, 88] represents an emerging and promising area of research that combines the strengths of retrieval-based methods with code generation techniques. Hayati et al.[82] introduced RECODE, the first model to integrate information retrieval into code generation. Parvez et al.Ā [87] proposed the REDCODER, which utilizes retrieval to enhance the input of the generation model, aiming to maximize the utilization of additional information. Drain et al.Ā [86] retrieve template functions from GitHub and relevant answers from Stack Overflow to improve code generation. Zhou et al.Ā [88] proposed Docprompting, which retrieve relevant APIs from documents and add them into prompts to improve code generation. Zan et al.Ā [89] proposed CodeGenAPI, which retrieve private APIs from documents to improve code generation. The APIRetriever module of CodeGenAPI and our approach both focus on API recommendation. However, CodeGenAPI directly retrieves APIs from documentation and cannot effectively handle high-level tasks that require combinations of APIs.
@@ VIII Conclusion and Future Work
Code generation within code repositories often necessitates comprehensive knowledge of the entire repository context. Existing techniques often lack such a holistic understanding of code repository knowledge, which is crucial for effective code generation. In this paper, we introduce the $A^{3}$-CodGen framework, which is designed to construct, mine, and unify three essential types of code repository knowledge: local-aware, global-aware, and third-party-library-aware knowledge. To validate the effectiveness of these knowledge types, we conducted benchmark experiments and identified optimal configurations for their utilization. The results demonstrate that the $A^{3}$-CodGen framework significantly enhances the ability of LLMs to fully utilize local, global, and third-party library information. This results in the generation of concise, high-quality code. In the future, we plan to expand the applicability of our framework to multiple programming languages. Furthermore, we aim to enhance our knowledge utilization methods to enable LLMs to become more aware of various types of code knowledge within code repositories.